OpenStudy (anonymous):

James has decided to start saving for a cruise over the summer. His money is currently in the local bank, modeled by the function s(x) = 102. He is able to do work around the neighborhood to earn extra money and that function would be a(x) = 8(x + 2), where x is measured in hours. Explain to James how he can create a function that combines the two, and describe any simplification that can be done. @Asevilla5 @IMStuck @aum @math&ing001

OpenStudy (anonymous):

Please explain

OpenStudy (math&ing001):

Simply add the two functions s(x)+a(x) then simplify.

OpenStudy (anonymous):

show and explain with words please

OpenStudy (aum):

Let m(x) be the total money function. x is in hours m(x) = s(x) + a(x) = 102 + 8(x+2) = 102 + 8x + 16 = 8x + 118

OpenStudy (anonymous):

okay

OpenStudy (asevilla5):

I'd go with s(x)+a(x) which is 102+8(x+2)

OpenStudy (aum):

The problem states: "He is able to do work around the neighborhood to earn extra money and that function would be a(x) = 8(x + 2), where x is measured in hours." We add them because the problem states: "create a function that combines the two". s(x) + a(x) = 102 + 8(x+2) = 8x + 118